OpenStudy (anonymous):

solve the system of equations using substitution y=3x-1 5x+2y=9 a.(1,2) b.(3,1) c.(2,1) d.(1,3)

OpenStudy (mathmate):

Substitution means that you would replace a variable used in one equation by the definition in another. Example here: y=3x-1 defines (explicitly) what y is. We can substitute y by 3x-1 in the second equation to get: 5x+2(3x-1)=9 which is one equation in one unknown. From here, we can solve for x, and subsequently for y.

OpenStudy (butterflydreamer):

the simple explanation LOL. When a question asks you to use substitution.. we can see from the first line that they ALREADY give you what "y" equals to (in this case y=3x-1) So in the second equation it says 5x+2y = 9 And since we know that y=3x-1 , we would substitute the value of y into the second equation. 5x+ 2(3x-1) =9 Then we expand. 5x + 6x-2 = 9 Collect like terms and solve for x. 11x - 2 = 9 11x= 11 x= 1 Then since we know what "x" equals to, we substitute that value into y=3x-1 Sooooo, y=3(1) -1 y=3-1 y=2. So your answer is (1,2) and thus "a"

OpenStudy (anonymous):

thanks

OpenStudy (butterflydreamer):

you are welcome :)